# Automated Synthesis of Formally Verified Multi-Abstraction Function Summaries

### v1

@@ III-B Symbolic Execution
Denote the function’s exit point by $p_{-1}$, then $\mathrm{SymExecCollect}(\mathcal{F},\textit{Pre},p_{-1})$ provides all accumulated path conditions and states at termination of $\mathcal{F}$.
@@ VI-B Implementation Details
A key feature of our toolchain is the interoperability between VST-A and Frama-C, despite their different assertion languages. During the $\mathrm{LoopInvGen}$ phase, LLM-generated loop invariants in ACSL are translated into VST user assertion format. This enables VST-A to correctly interpret and leverage them for symbolic execution across loop and function boundaries. After symbolic execution, VST-A’s generated RSPs are converted from its assertion format to ACSL for subsequent Frama-C automated verification. To achieve these, we establish an automatic translation that enables the conversion of assertions in both directions (see the supplementary material).
@@ VII-A1 Performance on benchmarks
syGas GPT-3.5-turbo 94.74% 82.71% 69.92% 100.00% 93.98% 82.71% 100.00% 96.24% 87.22%
syGas GPT-4O-mini 89.37% 78.95% 65.41% 100.00% 91.73% 81.20% 100.00% 95.49% 85.71%
syGas GPT-4O 98.50% 93.98% 87.22% 100.00% 96.99% 94.74% 100.00% 99.25% 96.99%
syGas DeepSeek-V3 98.50% 94.74% 87.97% 100.00% 99.25% 98.50% 100.00% 99.25% 98.50%
syGas Claude-3.7-sonnet 100.00% 99.25% 97.74% 100.00% 100.00% 99.25% 100.00% 100.00% 99.25%
@@
Table I lists the performance results of our approach ARSPG on the three benchmarks. ARSPG notably excels at handling single-loop numerical programs (as seen in sysGus), and can also generate accurate invariants for multi-loop programs, despite their complexity. Moreover, these results show that both the use of stronger LLM models and consistent improvement across Pass@k significantly benefit performance on these more challenging problems.
@@ Discussion
We notice that the accuracy on OOPSLA and SV-COMP benchmarks declined compared to sysGus. This is due to two reasons:
@@ VII-A2 Comparison with existing work
Method syGas OOSPLA SV-COMP
@@
In Table II, we evaluated ARSPG under Pass@1 and zero-shot conditions against a range of existing tools for loop invariant generation, broadly categorized into traditional and LLM-based approaches. Traditional tools like CODE2INV[10], CLN2INV[11], and HOLA[8] are typically optimized for specific, single datasets. LLM-based tools such as AutoSpec and ACInv provided a benchmark, for which we compared to their best reported results.
ARSPG was assessed in two configurations: ARSPG, using the GPT-4O model with a Pass@5 setting, and ARSPG-Best, representing optimal performance with the Claude-3.7-sonnet model at Pass@5. As presented in Table II, when ARSPG is augmented with advanced LLMs, it outperforms CODE2INV and CLN2INV on the syGus benchmark, with ARSPG-Best achieving a remarkable 99.3%. On the more challenging OOPSLA benchmark, ARSPG-Best also demonstrates strong performance at 87.0%, outperforming both AutoSpec and ACInv. For the SV-COMP benchmark, our results are consistent with those of AutoSpec and ACInv, achieving 76.2%. These findings collectively underscore the effectiveness of our approach in generating high-quality loop invariants across diverse benchmarks.
@@ VII-C Summary Generation and Generalization on Motivating Example
First of all, we apply our tool to derive the RSP for TripleAbsMaxFun, the disjunction of formulas covering all 64 execution paths and capturing the function’s complete behavior. For clarity and page limit, the full summary of the function is presented in supplemental material.

### v2

@@ III-B Symbolic Execution
Denote the function’s exit point by $p_{-1}$, then $\mathrm{SECollect}(\mathcal{F},\textit{Pre},p_{-1})$ provides all accumulated path conditions and states at termination of $\mathcal{F}$.
@@ VI-B Implementation Details
A key feature of our toolchain is the interoperability between VST-A and Frama-C, despite their different assertion languages. During the $\mathrm{LoopInvGen}$ phase, LLM-generated loop invariants in ACSL are translated into VST user assertion format. This enables VST-A to correctly interpret and leverage them for symbolic execution across loop and function boundaries. After symbolic execution, VST-A’s generated RSPs are converted from its assertion format to ACSL for subsequent Frama-C automated verification. To achieve these, we establish an automatic translation that enables the conversion of assertions in both directions (see Appendix -B).
@@ VII-A1 Performance on benchmarks
syGus GPT-3.5-turbo 94.74% 82.71% 69.92% 100.00% 93.98% 82.71% 100.00% 96.24% 87.22%
syGus GPT-4O-mini 89.37% 78.95% 65.41% 100.00% 91.73% 81.20% 100.00% 95.49% 85.71%
syGus GPT-4O 98.50% 93.98% 87.22% 100.00% 96.99% 94.74% 100.00% 99.25% 96.99%
syGus DeepSeek-V3 98.50% 94.74% 87.97% 100.00% 99.25% 98.50% 100.00% 99.25% 98.50%
syGus Claude-3.7-sonnet 100.00% 99.25% 97.74% 100.00% 100.00% 99.25% 100.00% 100.00% 99.25%
@@
Table I lists the performance results of our approach ARSPG on the three benchmarks. ARSPG notably excels at handling single-loop numerical programs (as seen in syGus), and can also generate accurate invariants for multi-loop programs, despite their complexity. Moreover, these results show that both the use of stronger LLM models and consistent improvement across Pass@k significantly benefit performance on these more challenging problems.
@@ Discussion
We notice that the accuracy on OOPSLA and SV-COMP benchmarks declined compared to syGus. This is due to two reasons:
@@ VII-A2 Comparison with existing work
Method syGus OOSPLA SV-COMP
@@
In Table II, we evaluated ARSPG under Pass@5 and zero-shot conditions against a range of existing tools for loop invariant generation, broadly categorized into traditional and LLM-based approaches. Traditional tools like CODE2INV[10], CLN2INV[11], and HOLA[8] are typically optimized for specific, single datasets. LLM-based tools such as AutoSpec and ACInv provided a benchmark, for which we compared to their best reported results.
ARSPG was assessed in two configurations: ARSPG, using the GPT-4O, and ARSPG-Best, representing optimal performance with the Claude-3.7-sonnet. As presented in Table II, when ARSPG is augmented with advanced LLMs, it outperforms CODE2INV and CLN2INV on the syGus benchmark, with ARSPG-Best achieving a remarkable 99.3%. On the more challenging OOPSLA benchmark, ARSPG-Best also demonstrates strong performance at 87.0%, outperforming both AutoSpec and ACInv. For the SV-COMP benchmark, our results are consistent with those of AutoSpec and ACInv, achieving 76.2%. These findings collectively underscore the effectiveness of our approach in generating high-quality loop invariants across diverse benchmarks.
@@ VII-C Summary Generation and Generalization on Motivating Example
First of all, we apply our tool to derive the RSP for TripleAbsMaxFun, the disjunction of formulas covering all 64 execution paths and capturing the function’s complete behavior. For clarity and page limit, the full summary of the function is presented in Appendix -C